# version 1.0

### v6

@@ 11.2.1 Generalization to High Dimension
Consider a linear dynamical system with matrices $A,B,C$ and further assume that $A$ is symmetric. This implies that $A$ can be diagonalized over the set of reals as $A=U^{\top}DU$. Decomposing over each dimension $j\in[d]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
$\mathbf{y}_{t}-\mathbf{y}_{t-1}-CB\mathbf{u}_{t-1}$
$=\sum_{i=1}^{t-1}CU^{\top}(D^{i-1}-D^{i-2})UB\mathbf{u}_{t-1-i}$
$=\sum_{i=1}^{t-1}CU^{\top}(\sum_{j=1}^{d_{x}}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})e_{j}e_{j}^{\top})UB\mathbf{u}_{t-i-1}$
$=\sum_{j=1}^{d_{x}}c_{j}b_{j}^{\top}\sum_{i=1}^{t-1}(\alpha^{i-1}_{j}-\alpha_{j}^{i-2})\mathbf{u}_{t-1-i}$ $c_{j},b_{j}\leftarrow j^{th}\mbox{ row/col of }CU^{\top},UB$
$=\sum_{j=0}^{d_{x}}\left(M^{\star}_{j}\mathbf{u}_{t}+{M}^{\star}_{j}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})\right)$
$=\tilde{M}^{\star}\mathbf{u}_{t}+\tilde{M}^{\star}\mu_{\alpha_{j}}(\mathbf{u}_{1:t-1})$
Thus, we can apply the same approximation techniques from the previous section, and derive an analogous conclusion to Lemma 11.3 for dimension higher than one. Crucially, the number of parameters required does not depend on the hidden dimension of the original system. The details are left as Exercise 11.6.
@@ 11.3 Online Spectral Filtering
The prediction class $\prod_{h}^{\lambda}$ can be learned by online gradient descent (Algorithm 3) in the online convex optimization framework, which gives rise to the following algorithm.
Figure 11.2: The filters obtained by the eigenvectors of $Z$.
Algorithm 7 Online Spectral Filtering
@@
2: Compute $h$ top eigenvectors of $Z_{T}$, denoted $\phi_{1},...,\phi_{h}$.
@@
4:  Predict $\hat{\mathbf{y}}_{t}=M_{0}^{t}\mathbf{u}_{t}+\sum_{j=1}^{h}M^{t}_{j}\phi_{j}(\tilde{\mathbf{u}}_{t-i})$ , as per definition 11.2, and observe true $\mathbf{y}_{t}$.
5:  Define cost function $f_{t}(M)=\|\mathbf{y}_{t}-\tilde{\mathbf{y}}_{t}\|^{2}$.
@@
The eigenvectors of $Z$ have an interesting structure: their magnitude does not decay with time, which explains their improved approximation as captured formally in Lemma 11.3.
The regret guarantees of the online gradient descent algorithm, as given in Theorem 6.1, directly imply the following corollary.
Theorem 11.5.
$\sum_{t}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\prod_{h}^{\lambda}}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}\leq 2GD\sqrt{T}.$
@@ 11.4 Conclusion
Looking ahead, Spectral Filtering opens the door for several exciting directions. Extensions to asymmetric system matrices, as referenced in the bibliographic section, provide a pathway toward broader applicability. Moreover, integrating spectral methods with control strategies may lead to novel approaches for nonstochastic control with partial observations, which we explore in the next chapter.
@@ 11.5 Bibliographic Remarks
Deviating from classical recovery and into the world of improper prediction, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel matrix $Z$. The methods we surveyed in previous chapters to learning linear dynamical systems had statistical and computational complexity that depended on the hidden dimension of the system. In contrast, the spectral filtering technique studied in this chapter has regret bounds and computational complexity guarantees that do not depend on the hidden dimension.
The spectral properties of the Hankel matrices used in this chapter are from [BT17].
The first extension of the spectral learning technique to asymmetric matrices is due to [HLS+18]. A more recent advancement of spectral filtering for general linear dynamical systems was recently put forth in [MH25].
The technique of spectral filtering has found applications in control and sequence prediction since its introduction.
In [AHL+18], the authors use a spectral filtering technique to concisely represent linear dynamical systems in a linear basis, formulating the optimal control problem as a convex program. This method circumvents the nonconvex challenges associated with explicit system identification and latent state estimation, providing provable optimality guarantees for the control signal. Closed-loop control using spectral filtering techniques was recently advanced in [BBDH25].
A recent line of work makes use of the spectral basis discovered in the line of work surveyed in this chapter to design neural architectures for sequence prediction. This was first suggested in the context of state space models in the spectral basis in [ASCH23], and was extended to hybrid models in [LND+24]. Later extensions explored efficient implementation [ACD+24] and the concept of length generalization in sequence prediction [MDA+24].
@@ 11.6 Exercises
11.1. Generalize Lemma 11.3 to all real numbers in the range [1,1][-1,1], instead of $[0,1]$.
11.2. Generalize Lemma 11.3 to any dimension $d_{x}>1$.

### v8

@@ 11.2.1 Generalization to High Dimension
Consider a linear dynamical system with matrices $A,B,C$. The expression $\tilde{\mathbf{u}}_{t}$ is now a block matrix of dimensions $T\times d_{u}$, and thus $\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha}$ is a vector of dimension $d_{u}$.
We further assume that $A$ is symmetric. This implies that $A$ can be diagonalized over the set of reals as $A=U^{\top}DU$. Decomposing over each dimension $j\in[d_{x}]$ separately, and denoting $D_{jj}=\alpha_{j}$, we can write (11.1):
$\mathbf{y}_{t}=\sum_{i=1}^{t-1}CU^{\top}D^{i-1}UB\mathbf{u}_{t-i}$
$=\sum_{i=1}^{t-1}CU^{\top}\left(\sum_{j=1}^{d_{x}}\alpha^{i-1}_{j}e_{j}e_{j}^{\top}\right)UB\mathbf{u}_{t-i}$
$=\sum_{j=1}^{d_{x}}c_{j}b_{j}^{\top}\sum_{i=1}^{t-1}\alpha^{i-1}_{j}\mathbf{u}_{t-i}$ $c_{j},b_{j}\leftarrow j^{th}\mbox{ col of }CU^{\top},UB$
$=\sum_{j=1}^{d_{x}}M_{j}\big(\tilde{\mathbf{u}}_{t}^{\top}\mu_{\alpha_{j}}\big)$ $\text{where }M_{j}:=c_{j}b_{j}^{\top}\in{\mathbb{R}}^{d_{y}\times d_{u}}.$
Thus, we can apply the same approximation techniques from the previous section, and derive an analogous conclusion to Theorem 11.3 for dimension higher than one. Crucially, the number of parameters required does not depend on the hidden dimension of the original system. The details are left as Exercise 11.6.
@@ 11.3 Online Spectral Filtering
The prediction class $\Pi_{h}$ can be learned by online gradient descent in the online convex optimization framework, spelled out in Algorithm 7. The regret guarantees of the resulting algorithm, as given in Theorem 6.1, directly imply the following.
Theorem 11.5.
$\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi_{h}}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\mathbf{y}_{t}^{\pi}\|^{2}\leq 2GD\sqrt{T}.$
Algorithm 7 Online Spectral Filtering with OGD
@@
2: Compute $h$ top eigenvectors of $Z_{T}$, denoted $\phi_{1},\dots,\phi_{h}$.
@@
4:  Predict $\hat{\mathbf{y}}_{t}=\sum_{j=1}^{h}M^{t}_{j}\big(\tilde{\mathbf{u}}_{t}^{\top}\phi_{j}\big)$ and observe true $\mathbf{y}_{t}$.
5:  Define loss $f_{t}(M)=\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}$.
@@
However, the gradients of the loss can depend polynomially on $T$, since the feature vectors can have norms that grow with the horizon. This renders the regret guarantee of OGD uninformative in our setting.
To address this issue, we turn instead to the Vovk-Azoury-Warmuth (VAW) forecaster from Appendix A. In the spectral filtering setting, the feature vector is known before the prediction is made, so the VAW protocol is a natural fit. For simplicity of exposition, we present the scalar setting ($d_{y}=d_{u}=1$). The resulting regret depends multiplicatively only on the label bound $Y$, while the dependence on the feature magnitudes appears only through a logarithmic determinant term.
Algorithm 8 Spectral Filtering with VAW
1:Input: Filter size $h$, $\lambda>0$, horizon $T$.
2:Precompute: Top $h$ eigenvectors of $Z_{T}$, stored in $\Phi_{h}\in{\mathbb{R}}^{T\times h}$.
3:Initialize: $A_{0}=\lambda I_{h}$, $v_{0}=0\in{\mathbb{R}}^{h}$.
4:for $t=1,2,\dots,T$ do
5:  $\tilde{\mathbf{u}}_{t}=[u_{t-1},\dots,u_{1},0,\dots,0]^{\top}$.
6:  $h_{t}=\Phi_{h}^{\top}\tilde{\mathbf{u}}_{t}$.
7:  Curvature update: $A_{t}=A_{t-1}+h_{t}h_{t}^{\top}$.
8:  Predict $\hat{y}_{t}=h_{t}^{\top}A_{t}^{-1}v_{t-1}$ and observe $y_{t}$.
9:  Update target vector: $v_{t}=v_{t-1}+y_{t}h_{t}$.
10:end for
Corollary 11.6 (VAW regret for spectral filtering).
Assume $|y_{t}|\leq Y$ for all $t$, and let $\Theta\subseteq{\mathbb{R}}^{h}$ satisfy $\|\theta\|_{2}\leq B$ for all $\theta\in\Theta$. Then
$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\min_{\theta\in\Theta}\sum_{t=1}^{T}(y_{t}-\theta^{\top}h_{t})^{2}\;\leq\;\lambda B^{2}+Y^{2}\log\det\!\left(I_{h}+\frac{1}{\lambda}\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right).$
If in addition $|u_{s}|\leq 1$ for all $s$, and setting $\lambda=1/B^{2}$, gives
$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\min_{\theta\in\Theta}\sum_{t=1}^{T}(y_{t}-\theta^{\top}h_{t})^{2}\;\leq\;1+hY^{2}\log\!\left(1+B^{2}T^{2}\right).$
Indeed, applying the VAW regret bound from Appendix A with $a_{t}=h_{t}$ and $b_{t}=y_{t}$ yields
$\sum_{t=1}^{T}(y_{t}-\hat{y}_{t})^{2}-\sum_{t=1}^{T}(y_{t}-\theta^{\top}h_{t})^{2}\;\leq\;\lambda\|\theta\|_{2}^{2}+Y^{2}\log\det\!\left(I_{h}+\frac{1}{\lambda}\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right)$
for every $\theta\in{\mathbb{R}}^{h}$. Taking $\theta$ to be the best predictor in $\Theta$ gives the corollary.
Since $\Phi_{h}$ has orthonormal columns, $\|\Phi_{h}^{\top}\|_{2}=1$. If $|u_{s}|\leq 1$, then $\|\tilde{\mathbf{u}}_{t}\|_{2}\leq\sqrt{T}$ and hence $\|h_{t}\|_{2}\leq\sqrt{T}$. Therefore $\|h_{t}h_{t}^{\top}\|_{2}\leq T$, so
$\left\|\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right\|_{2}\leq T^{2},$
which implies
$\log\det\!\left(I_{h}+\frac{1}{\lambda}\sum_{t=1}^{T}h_{t}h_{t}^{\top}\right)\leq h\log\!\left(1+\frac{T^{2}}{\lambda}\right).$
Even if the feature magnitudes grow polynomially with $T$, their contribution remains only logarithmic. Thus VAW avoids the large multiplicative constants that arise in first-order methods, and the leading multiplicative dependence is only on the label magnitude $Y$.
@@ 11.4 Conclusion
Looking ahead, Spectral Filtering opens the door for several exciting directions. Extensions to asymmetric system matrices, as referenced in the bibliographic section, provide a pathway toward broader applicability. Moreover, integrating spectral methods with control strategies may lead to novel approaches for nonstochastic control with partial observations, which we explore in the next part of this text.
@@ 11.5 Bibliographic Remarks
Deviating from classical recovery and into the world of improper prediction, the work of [HSZ17] devised an improper learning technique using the eigenvectors of the Hankel moment matrix $Z$. The methods we surveyed in previous chapters for learning linear dynamical systems had statistical and computational complexity that depended on the hidden dimension. In contrast, the spectral filtering technique studied in this chapter has regret bounds and computational complexity guarantees that do not depend on the hidden dimension.
The spectral decay properties of the Hankel (Hilbert) matrices used in this chapter can be derived from the displacement-structure bounds of [BT17], and are closely connected to the classical approximation theory of Hankel operators; see, e.g., [PEL03] for a comprehensive treatment.
Connections to realization and model-reduction theory trace back to the Ho–Kalman construction [HK66] and to balanced truncation and related approximation methods [ANT05].
Spectral filtering can also be interpreted through the lens of operator approximation: it approximates the input–output map via projection onto leading modes of a compact Hankel operator. This viewpoint is closely related to Koopman/operator-theoretic approaches and data-driven spectral methods such as Dynamic Mode Decomposition [TRL+14].
The first extension of spectral filtering beyond symmetric settings to general (asymmetric) linear dynamical systems is due to [HLS+18]. More recent dimension-free advances for learning asymmetric linear dynamical systems were developed in [MH25].
The technique of spectral filtering has found applications in control and sequence prediction since its introduction. In [AHL+18], the authors use a spectral filtering representation to cast control of unknown linear dynamical systems as a convex program. Closed-loop control using spectral filtering techniques was recently advanced in [BBD+25].
A recent line of work makes use of spectral bases to design neural architectures for sequence prediction. This was first suggested in the context of spectral state space models [ASC+23], and was extended to hybrid models such as FlashSTU [LND+24]. Later extensions explored efficient implementation [ACD+24] and the concept of length generalization [MDA+24]. Structured state-space models such as S4 [GGR22] also exploit diagonalizations of linear recurrences, though derived from continuous-time parameterizations rather than Hankel moment structure.
At a conceptual level, spectral filtering replaces explicit state estimation by operator approximation: instead of reconstructing latent states, it approximates the input–output map in a near-optimal spectral basis. This perspective unifies ideas from realization theory, compact operator approximation, and modern sequence modeling.
@@ 11.6 Exercises
11.1. Generalize Theorem 11.3 to all real numbers in the range [1,1][-1,1], instead of $[0,1]$.
11.2. Generalize Theorem 11.3 to any dimension $d_{x}>1$.